# Phase space noncommutativity, power-law inflation and quantum cosmology

### v2

@@ 1 Introduction
Models based on noncommutative (NC) theories are still of interest. The main motivation for employing noncommutativity of the canonical type may be traced back to the investigations that establish its relationship with string and M theory [1]. In addition, the works in semi–classical gravity [2] could be related to NC field theories. Furthermore, the application of NC theories allows for the consideration of several fascinating aspects, such as physics at extremely small distances [3], non-locality, IR/UV mixing [4], Lorentz violation [5], the NC Landau problem [6] and the quantum Hall effect [7].
@@ 2 Noncommutative SB cosmology
From using the standard relationship between the Lagrangian and the classical Hamiltonian $\mathcal{H}$, it is straightforward to show that

### v3

@@ 1 Introduction
Models based on noncommutative (NC) theories are still of interest. The main motivation for employing noncommutativity of the canonical type may be traced back to the investigations that establish its relationship with string and M theory [1]. In addition, the works in semi–classical gravity [2] could be related to NC field theories. Furthermore, the application of NC theories allows for the consideration of several fascinating aspects, such as physics at extremely small distances [3], non-locality, IR/UV mixing [4], Lorentz violation [5], the NC Landau problem [6] and the quantum Hall effect [7]; as other interesting applications of the noncommutativity in physics, see also [8, 9, 10].
@@ 2 Noncommutative SB cosmology
$m=4$ From using the standard relationship between the Lagrangian and the classical Hamiltonian $\mathcal{H}$, it is straightforward to show that
@@ 7 Summary and discussions
Let us conclude this section by underlying a few essential points.
  • 1.
    It is noteworthy that, in general, one might not consider the proper transformations that provide the corresponding usual gravitational model with the canonical kinetic term from the action (2). Nonetheless, given the significance of the subject, let us consider a specific transformation $d\tilde{\phi}=\sqrt{J(\phi)}d\phi$ that takes the gravitational sector of the SB model (2) to [43] (see also [44])
    ${\cal S}^{{}^{(D)}}=\int d^{{}^{\,D}}\!x\sqrt{-g}\,\Big{[}R^{{}^{(D)}}-g^{\alpha\beta}\,({\nabla}_{\alpha}\tilde{\phi})({\nabla}_{\beta}\tilde{\phi})-U(\tilde{\phi})\Big{]},$ (86)
    where ${\cal W}\phi^{n}\equiv J(\phi)>0$ and $U[\tilde{\phi}(\phi)]=V(\phi)$.
    Moreover, by expressing $J(\phi)$ in terms of the potentials, equation (86) transforms to
    ${\cal S}^{{}^{(D)}}=\int d^{{}^{\,D}}\!x\sqrt{-g}\,\Big{[}R^{{}^{(D)}}-\left(V_{,\phi}\frac{dU^{-1}(V(\phi))}{dV(\phi)}\right)^{2}g^{\alpha\beta}\,({\nabla}_{\alpha}\phi)({\nabla}_{\beta}\phi)-V({\phi})\Big{]},$ (87)
    where $U^{-1}$ denotes the inverse function of $U$.
    We observe that (86) and (87) are equivalent actions that could provide identical predictions [43]. It is crucial to emphasize that for any non-canonical model with the specific coupling function (see, for example, the SB model with $J(\phi)={\cal W}\phi^{n}>0$) using the previously mentioned transformation to produce the canonical kinetic term may restrict one to take a particular canonical potential. Furthermore, in actual practice, when a particular cosmological model is studied, it is sometimes much simpler to work with only one of the equivalent actions (i.e. either (2) or (86)) than the other. This makes it easier to obtain analytical or numerical results and to analyze the consequences. With the other equivalent action, however, it can be more challenging or even impossible.
    It is important to highlight that, the NC models we presented here, most notably the NC quantum model (see section 6), we believe that, to the best of our knowledge, the work we have done is completely new and has nothing in common with earlier models, even when the transformations mentioned above are taken into consideration. Moreover, in regard to subsection 4.2, we stress that our NC classical model, which is based on the simple deformed phase space algebra, produces the same results as the intriguing NC inflationary model found in Ref. [28]. This, in turn, indicates a correspondence between these gravitational models and can be applied in a variety of contexts for both model parameters comparison and structural phenomenological discussions. Furthermore, subsection 4.1 also represents interesting numerical consequences. Concretely, it not only indicates an inflationary model that can solve the problems associated with the equivalent commutative models, but also corresponds with the Starobinsky model, see for instance [37], and references therein.
    Eventually, we believe that the models developed in this paper, which consider an intriguing generalized noncommutativity, arbitrary dimensions, and a gravitational scalar model with a non-canonical kinetic term including additional parameters, have yielded new and interesting results, and have the potential to represent correspondence with previous quantum gravity models.
  • 2.